the length of a rectangle is twice its width. If the peremiter of the rectangle is 48ft, find its area
Answer by Maths68(1474)

Perimeter of rectangle = P = 48ft
Length = 2w
Width = w

Perimeter of rectangle = 2(Length+Width)
48=2(2w+w)
48=2(3w)
48=6w
48/6=6w/6
8=w
w=8ft

Length = 2w = 2(8)=16ft
Width = w = 8ft

Area of rectangle = Length * Width
Area of rectangle = 8 * 16
Area of rectangle = 128 square feet.